# Metrology of microwave fields based on trap-loss spectroscopy with cold Rydberg atoms

### v1

@@ III Discussion
IV Methods
IV.1 Experimental setup details
We create a 87Rb MOT in a glass cell with a 26 G/cm magnetic field gradient and 3 pairs of retro-reflected beams at 780 nm with 1/e2 diameter around 9 mm, used for both cooling and repumping. The dimensions of the atomic cloud are typically on the order of 1 mm. The cooling beams come from a single Ortel telecom laser diode at 1560 nm, frequency-doubled with a PPLN waveguide crystal. The frequency of the cooling light is detuned by 20 MHz from the |$\ket{5S_{1/2},F=2}\leftrightarrow\ket{5P_{3/2},F^{\prime}=3}$ transition. The repumping light comes from the phase modulation of the cooling light, with a power ratio beetween the repumping and cooling frequencies around 20 %. The MOT beams carry a total amount of power around 180 mW. The coupling between the ground state and the |$\ket{61S_{1/2}}$ Rydberg state is performed with a 2 photon scheme (780 nm and 480 nm), with about 500 MHz detuning from the intermediate |$\ket{5P_{3/2},F^{\prime}=3}$ state. Both beams are linearly polarized, their direction of polarization being adjusted to maximize the contrast of the spectral feature shown in figure 2(a). The two counterpropagating Rydberg laser beams have a $1/e^{2}$ diameter of 1 mm, and their respective power are 2 mW for the infrared one (frequency-doubled Rio Planex telecom laser diode at 1560 nm) and 80 mW for the blue one (Moglabs injection-locked ILA system). The two lasers are frequency-stabilized using a Pound-Drever-Hall lock on an ultrastable cavity from SLS with a finesse of the order of 2100 for the infrared light and 18000 for the blue one. The sweep of the two-photon detuning is achieved by locking a sideband of the phase-modulated infrared Rybderg laser to the reference cavity while sending the carrier to the atoms, and then sweeping the frequency of the phase modulation. The MW signal sent onto the atoms is provided by a Rohde & Schwarz SMB100A generator stabilized by a 10 MHz reference from a GPS receiver. The horn antenna is set 10 cm away from the atoms.
IV.2 Theoretical model and rate equations
@@
IV.3 Autler Townes effect caused by the cooling light
As discussed in section II.1, a close look at the experimental curve of figure 2(a) reveals the presence of a second spectral feature around $-35$ MHz. This feature can be explained by the cooling light operating on the |$\ket{5S_{1/2},F=2}\leftrightarrow\ket{5P_{3/2},F^{\prime}=3}$ transition with a Rabi frequency $\Omega_{MOT}=2\pi\times 29\,\textrm{MHz}$ and a detuning $\delta_{MOT}=2\pi\times-20\,\textrm{MHz}$, which results in an Autler-Townes splitting of the ground state also involved in the trap loss measurement. This is quantitatively confirmed by a numerical model based on Bloch equations and involving the three atomic states |$\ket{g}\equiv\ket{5S_{1/2},F=2}$, |$\ket{i}\equiv\ket{5P_{3/2},F^{\prime}=3}$ and |$\ket{r}\equiv\ket{61S_{1/2}}$, and two Rabi couplings: $\Omega_{MOT}$ between |$\ket{g}$ and |$\ket{i}$, and $\Omega_{2ph}=2\pi\times 9\;kHz$ between |$\ket{g}$ and |$\ket{r}$. By furthermore taking into account the decay rates $\Gamma_{\ket{i}}=2\pi\times 6.06\times 10^{6}\,s^{-1}$ and $\Gamma_{\ket{r}}=7.1\times 10^{3}\,s^{-1}$, and by introducing empirical coherence damping rates $\gamma_{gi}=2\pi\times 2$ MHz and $\gamma_{gr}=2\pi\times 4$ MHz, we obtain the red trace on figure 6 which shows a good agreement with the measured data (identical to the one presented on figure 2(a)).
@@ IV.4 Influence of the MOT dynamics
We report on figure 8 the estimated frequency $\omega_{dip}/2\pi$ corresponding to the minimum fluorescence for each trace, as a function of the scanning rate and direction. As expected the estimation error increases with the scanning rate, but the average between the estimated frequencies using an ascending and a descending scan remains stable at the MHz level. The estimated frequencies based on the theoretical traces, also shown on figure 8, are in good agreement with the experimental points. Interestingly, the value of $\omega_{dip}/2\pi$ predicted by the model in the limit of very slow scanning rates is not zero but $\simeq 200$ kHz. This is due to the fact that the Rabi frequency $\Omega_{2ph}$ of the effective two-photon coupling between the ground and Rydberg states has a dependence in $\delta$ (see section IV.2), breaking the symmetry between ascending and descending scans.
@@
IV.5 Additional detuning from the MW source
@@ Acknowledgement
Author Contributions
The experiments, data analysis and numerical simulations were carried out by R.D., A.Bo., R.G., Q.M., C.B. and S.S. All work was supervised by A.Bo. and S.S. All authors discussed the results and contributed to the manuscript.

### v2

@@ V Conclusion
VI Appendix
VI.1 Experimental setup details
We create a 87Rb MOT in a glass cell with a 26 G/cm magnetic field gradient and 3 pairs of retro-reflected beams at 780 nm with 1/e2 diameter around 9 mm, used for both cooling and repumping. We estimate the partial pressure of rubidium in the cell to be on the order of $10^{-8}$ Torr. The dimensions of the atomic cloud are typically on the order of 1 mm. The cooling beams come from a single Ortel telecom laser diode at 1560 nm, frequency-doubled with a PPLN waveguide crystal. The frequency of the cooling light is detuned by 20 MHz from the |$\ket{5S_{1/2},F=2}\leftrightarrow\ket{5P_{3/2},F^{\prime}=3}$ transition. The repumping light comes from the phase modulation of the cooling light, with a power ratio beetween the repumping and cooling frequencies around 20 %. The MOT beams carry a total amount of power around 180 mW. The coupling between the ground state and the |$\ket{61S_{1/2}}$ Rydberg state is performed with a 2 photon scheme (780 nm and 480 nm), with about 500 MHz detuning from the intermediate |$\ket{5P_{3/2},F^{\prime}=3}$ state. Both beams are linearly polarized, their direction of polarization being adjusted to maximize the contrast of the spectral feature shown in figure 2(a). The two counterpropagating Rydberg laser beams have a $1/e^{2}$ diameter of 1 mm, and their respective power are 2 mW for the infrared one (frequency-doubled Rio Planex telecom laser diode at 1560 nm) and 80 mW for the blue one (Moglabs injection-locked ILA system). The two lasers are frequency-stabilized using a Pound-Drever-Hall lock on an ultrastable cavity from SLS with a finesse of the order of 2100 for the infrared light and 18000 for the blue one. The sweep of the two-photon detuning is achieved by locking a sideband of the phase-modulated infrared Rybderg laser to the reference cavity while sending the carrier to the atoms, and then sweeping the frequency of the phase modulation. The MW signal sent onto the atoms is provided by a Rohde & Schwarz SMB100A generator stabilized by a 10 MHz reference from a GPS receiver. The horn antenna is set 10 cm away from the atoms.
VI.2 Theoretical model and rate equations
@@
VI.3 Autler Townes effect caused by the cooling light
As discussed in section II.1, a close look at the experimental curve of figure 2(a) reveals the presence of a second spectral feature around $-35$ MHz. This feature can be explained by the cooling light operating on the |$\ket{5S_{1/2},F=2}\leftrightarrow\ket{5P_{3/2},F^{\prime}=3}$ transition with a Rabi frequency $\Omega_{MOT}=2\pi\times 29\,\textrm{MHz}$ and a detuning $\delta_{MOT}=2\pi\times-20\,\textrm{MHz}$, which results in an Autler-Townes splitting of the ground state also involved in the trap loss measurement. This is quantitatively confirmed by a numerical model based on Bloch equations and involving the three atomic states |$\ket{g}\equiv\ket{5S_{1/2},F=2}$, |$\ket{i}\equiv\ket{5P_{3/2},F^{\prime}=3}$ and |$\ket{r}\equiv\ket{61S_{1/2}}$, and two Rabi couplings: $\Omega_{MOT}$ between |$\ket{g}$ and |$\ket{i}$, and $\Omega_{2ph}=2\pi\times 9\;\textrm{kHz}$ between |$\ket{g}$ and |$\ket{r}$. By furthermore taking into account the decay rates $\Gamma_{\ket{i}}=2\pi\times 6.06\times 10^{6}\,s^{-1}$ and $\Gamma_{\ket{r}}=7.1\times 10^{3}\,s^{-1}$, and by introducing empirical coherence damping rates $\gamma_{gi}=2\pi\times 2$ MHz and $\gamma_{gr}=2\pi\times 4$ MHz, we obtain the red trace on figure 6 which shows a good agreement with the measured data (identical to the one presented on figure 2(a)).
@@ VI.4 Influence of the MOT dynamics
We report on figure 8 the estimated frequency $\omega_{dip}/2\pi$ corresponding to the minimum fluorescence for each trace, as a function of the scanning rate and direction. As expected the estimation error increases with the scanning rate, but the average between the estimated frequencies using an ascending and a descending scan remains stable at the MHz level. The estimated frequencies based on the theoretical traces, also shown on figure 8, are in good agreement with the experimental points. Interestingly, the value of $\omega_{dip}/2\pi$ predicted by the model in the limit of very slow scanning rates is not zero but $\simeq 200$ kHz. This is due to the fact that the Rabi frequency $\Omega_{2ph}$ of the effective two-photon coupling between the ground and Rydberg states has a dependence in $\delta$ (see section VI.2), breaking the symmetry between ascending and descending scans.
@@
VI.5 Additional detuning from the MW source